1(b) A software company is offering to a smart phone app developer the following contract. At 8% interest, what is the present worth of this contract? S30,000/year for the next 5 years S35,000/year for another 5 years (i.e. 5 years following the initial 5 years).
Answer #1

Answer:

Given

interest rate r=8%

for 1-5 years Contract A =$30,000

for 6-10 years Contract B= $35,000

So present worth of contract =A*(1-(1+r)^-5)/r+B*(1-(1+r)^-5)/{r*(1+r)^5}

PW= 30000*(1-(1+8%)^-5)/8%+35000*(1-(1+8%)^-5)/{8%*(1+8%)^5}=$214889.298